OpenStudy (anonymous):

The grades received by 200 students follow a normal distribution. The mean of the grades is 70%, and the standard deviation is 7%. The number of students who received a grade greater than 70% is about (blank) and the number of students who got a grade higher than 84% is about (blank)

OpenStudy (anonymous):

I would love to answer this for you, but what do the "about (blank)s represent?

OpenStudy (anonymous):

\[P(X>70)=P\left(\frac{X-70}{7}>\frac{70-70}{7}\right)=P(Z>0)\] \[P(X>84)=P\left(\frac{X-70}{7}>\frac{84-70}{7}\right)=P(Z>2)\]